blog:20240413

20240413

1800-

Sam, Azzan, Roy, Derek

PHYS 401, 433C, Dome, Roof

  1. Review standard operating procedures
  2. Take calibration frames
  3. Take seeing frames
  4. Find 1 or 2 standard fields to observe
Scope Camera
Main ASI 432
Finder ASI 2600
Time Observed Source Name Filter Int Time [s] Gain Airmass Mag (filter) RA [h] DEC [deg] Source Type Scope Purpose Notes
20240413T1930 Flats U 12×0.1397 150 N/A N/A N/A N/A Flat Main Calibration Using SC 4.1 GARB
20240413T1930 Flats B 12×0.0327 150 N/A N/A N/A N/A Flat Main Calibration Using SC 4.1 GARB
20240413T1930 Flats V 12×0.0327 150 N/A N/A N/A N/A Flat Main Calibration Using SC 4.1 GARB
20240413T1930 Flats R 12×0.0327 150 N/A N/A N/A N/A Flat Main Calibration Using SC 4.1 GARB
20240413T1930 Flats I 12×0.0676 150 N/A N/A N/A N/A Flat Main Calibration Using SC 4.1 GARB
20240413T2000 Pollux V 150 N/A N/A N/A N/A Star Main Pointing Using SC 4.1
20240413T2030 HD 70272 V 3×0.25 0 1.011 4.25 8 43 Star Main Seeing
20240413T2030 HD 44537 V 3×0.5 0 1.162 4.9 6 49 Star Main Seeing Using SC 4.1
20240413T2030 HD 31314 V 3×5 0 1.496 8.1 4 40 Star Main Seeing Using SC 4.1
20240413T2030 HD 76351 V 3×1.5 0 1.135 5.4 8 11 Star Main Seeing Using SC 4.1
20240413T2145 SA 101 U 6×129 150 1.303 9-14 10 0 Landolt Field Main Data Redux SC 4.1, 15.011, ag-on
20240413T2145 SA 101 V 6×8 150 1.327 9-14 10 0 Landolt Field Main Data Redux SC 4.1, 15.011, ag-on
20240413T2145 SA 101 B 6×13 150 1.319 9-14 10 0 Landolt Field Main Data Redux SC 4.1, 15.011, ag-on
20240413T2145 SA 101 R 6×8 150 1.330 9-14 10 0 Landolt Field Main Data Redux SC 4.1, 15.011, ag-on
20240413T2145 SA 101 I 6×30 150 1.321 9-14 10 0 Landolt Field Main Data Redux SC 4.1, 15.011, ag-on
20240414T1630 Flat U 0.1×10 150 N/A N/A N/A N/A Flat Main Calibration
20240414T1630 Flat B 0.002×10 150 N/A N/A N/A N/A Flat Main Calibration
20240414T1630 Flat V 0.01×10 150 N/A N/A N/A N/A Flat Main Calibration
20240414T1630 Flat R 0.0005×10 150 N/A N/A N/A N/A Flat Main Calibration
20240414T1630 Flat I 0.01×10 150 N/A N/A N/A N/A Flat Main Calibration
20240414T1700 Dark N/A 129×10 150 N/A N/A N/A N/A Dark Main Calibration
20240414T1700 Dark N/A 30×10 150 N/A N/A N/A N/A Dark Main Calibration
20240414T1700 Dark N/A 13×10 150 N/A N/A N/A N/A Dark Main Calibration
20240414T1700 Dark N/A 8×10 150 N/A N/A N/A N/A Dark Main Calibration
20240414T1700 Dark N/A 0.1×10 150 N/A N/A N/A N/A Dark Main Calibration
20240414T1700 Dark N/A 0.002×10 150 N/A N/A N/A N/A Dark Main Calibration
20240414T1700 Dark N/A 0.001×10 150 N/A N/A N/A N/A Dark Main Calibration
20240414T1700 Dark N/A 0.0005×10 150 N/A N/A N/A N/A Dark Main Calibration
  1. Went through start-up procedures
  2. Moved dome so that sunlight wasn't heating-up dome
  3. Needed to reconnect data line for finder scope cam (ASI 2600)
  4. Took flat frames with sky at 1930, scope at zenith
    1. Unsure if 432 is north-aligned, OH WELL
    2. Forgot to record exposure times – will do this before taking darks, obv
  5. Finderscope camera disconnected (using SC 4.1)
    1. Azzan and Sam went up to reseat the data cable into the Icron
    2. The 2600 worked again, but then disconnected after maybe 60s
    3. Azzan and Sam then moved the 2600 data cable to LBD-2
    4. Seems to be working – FOR NOW
  6. Moved to Pollux for pointing – it was near zenith
  7. Dome tracking is working
  8. Pointing was off
    1. Got some starfield from the finderscope and uploaded to Astrometry.net (thanks to Connor for having an acct)
    2. Waited a few mins for astrometry.net to work
    3. Used the supplied RA and DEC to update the telescope position in DFM TCS
  9. Seems like focus might be off
  10. Went to HD 70272 to see if we can make things better
    1. Started at focus 1919.1
    2. Optimal focus is 1913.0
    3. FLATS ARE GARBAGE
    4. DELETING
  11. We Slew to HD 76351 the pointing seem to be off but we re-slew and it fixed the problem ( this is not expected)
  12. Started getting data for Dataredux study
    1. We started the autoguiding software phd2 (we re calibrate it)
    2. Source SA 101
  13. We took pretty pictures of different objects
  14. We did all the closing procedures (everything went smoothly)
  15. Took darks and flats next day after lecture ~1630-1730
    1. The sensor temperature seemed a bit too high for the darks, they'll be good for those flats, but we may want to retake darks for the SA 101 light frames

Seeing Targets:

- HD 70272, HD 44537, HD 31314, HD 76351

  1. Data cable for 2600 was disconnected – we reconnected to Icron
  2. Need to buy new casters for rolly ladder
  3. Need to get Azzan and Sam on Gold Card roof swipe access list
  4. Need to finish set
  5. 2600 seems to randomly disconnect
    1. Seems to be better behaved on LBD-2 for now?
    2. This needs to be investigated
Time Dome Temp Outside Temp Outside Humid
1915 18 20 23
2100